OpenStudy (anonymous):

1/2(2x-4)=-3/4(4x-16)

OpenStudy (anonymous):

distribute on both sides on the left, distribute the \(\frac{1}{2}\) to make \[\frac{1}{2}(2x-4)=x-2\]